OpenStudy (anonymous):

Solve the logarithmic equation for x. log x + log(x − 48) = 2

OpenStudy (kropot72):

Using the rules of logarithms \[\log_{} x+\log_{} (x-48)=\log_{}[x(x-48)]=2\] \[\log_{} 100=2\] \[x ^{2}-48x-100=0\] Now you need to solve the quadratic equation to find the values of x.